Calculate the work done by 2.4N force pushing a 400.g sandwich across a table 0.75m wide

1
Expert's answer
2022-02-18T08:22:21-0500

Given:

"F=2.4\:\rm N"

"\theta=0^\circ"

"d=0.75\:\rm m"


The work done

"W=Fd\cos\theta"

"W=2.4*0.75*\cos 0^\circ=1.8\:\rm J"